Use the cofactor expansion to determine

{{2000}{3120}{2-504}{1303}}


1
Expert's answer
2021-05-23T19:06:02-0400

"\begin{vmatrix}
 2 & 0&0&0 \\
 3 &1&2&0\\ 
2&-5&0&4\\
1&3&0&3\end{vmatrix}=2\begin{vmatrix}
 1 & 2&0 \\
 -5 & 0&4\\
3&0&3
\end{vmatrix}=2(-2)\begin{vmatrix}
 -5 & 4 \\
 3 & 3
\end{vmatrix}=-4(-15-12)=108."